# On the parameterized complexity of computing tree-partitions

### v4

@@ 5 Tree-cut width and the stability of tree-partition-width
In this section, we consider the relation of the notion of tree-cut width with (stable) tree-partition-width. Tree-cut width was introduced by Wollan [Wol15]. Ganian et al. [GKS15] showed that several problems that are $W[1]$-hard with treewidth as parameter are fixed parameter tractable with tree-cut width as parameter.
We begin by defining the tree-cut width of a graph $G=(V,E)$. A tree-cut decomposition $(T,\mathcal{X})$ consists of a rooted tree $T$ and a family $\mathcal{X}$ of bags $(X_{i})_{i\in V(T)}$ which form a near partition of $V(G)$ (i.e. some bags may be empty, but nonempty bags form a partition of $V(G)$). For $t\in V(T)$, we denote by $e(t)$ the edge of $T$ incident to $t$ and its parent. For $e\in E(T)$, let $T_{1},T_{2}$ denote the two connected components of $T-e$. We denote by $\operatorname{cut}(e)$ the set of edges with an endpoint in both of $\bigcup_{i\in V(T_{1})}X_{i}$ and $\bigcup_{i\in V(T_{2})}X_{i}$. The adhesion of $t\in V(T)$ is $\text{adh}(t)=|\operatorname{cut}(e(t))|$, and its torso-size is $\text{tor}(t)=|X_{t}|+b_{t}$ where $b_{t}$ is the number of edges $e\in E(T)$ incident to $t$ such that $|\operatorname{cut}(e)|\geq 3$. The width of the decomposition is then $\max_{t\in V(T)}\{\text{adh}(t),\text{tor}(t)\}$. Note that edges are allowed to go between vertices that are not in the same bag. The tree-cut width of a graph is the minimal width of tree-cut decomposition. When $|\operatorname{cut}(e)|\geq 3$, the edge $e$ is called bold, and otherwise, $e$ is called thin. When $\text{adh}(t)\leq 2$, node $t$ is called thin, otherwise it is called bold. In [GKS15], it is shown that a tree-cut decomposition can be assumed to be nice, meaning that if $t\in V(T)$ is thin then $N(Y_{t})\cap\left(\bigcup_{b\text{ sibling of }t}Y_{b}\right)=\varnothing$, where $Y_{i}$ is the union of $X_{j}$ for $j$ in the subtree of $i$.
Wollan [Wol15] shows that having bounded tree-cut width is equivalent to only having wall immersions of bounded size.
{observation}
@@
It is easy to see that $K_{3,n}$ has a tree-cut decomposition of width $O(\sqrt{n})$: we place $A$ as the center of a star, with about $\sqrt{n}$ leaves of size $\sqrt{n}$. Now we consider an arbitrary tree-cut decomposition of $K_{3,n}$ achieving width $O(\sqrt{n})$. We first note that the vertices of $A$ cannot be split into separate bags because if they were, any edge of the decomposition on the path between such bags would have adhesion at least $n/2$. Hence, there is a bag containing $A$ and we may root the tree of our decomposition in this bag. Each subtree of the decomposition will contribute to the torso-size, and each vertex will contribute linearly to the adhesion of the edge from its subtree to the root. Since we assume the width to be $O(\sqrt{n})$, we must have at most $O(\sqrt{n})$ vertices per subtree. Consequently, there must be $\Omega(\sqrt{n})$ subtrees, so the torso-size of the root is $\Omega(\sqrt{n})$.
@@
Proof 5.3.
Consider a nice tree-cut decomposition $(T,\mathcal{X})$ of a graph $G$ of width $k$. We will construct a tree-partition for a subdivision of $G$. Note that the bags are already disjoint, but some edges are not between neighbouring bags of $T$.
@@
Suppose that $u,v$ are in the same bag of $T$ and the edge $uv$ was subdivided to form the path $u,a_{1},\dots,a_{\ell},v$. We add the vertices $a_{i}$ in new bags containing, $\{a_{1},a_{\ell}\}$, $\{a_{2},a_{\ell-1},\dots$ which corresponds to a new branch of the decomposition of width at most $2$.
@@
This gives a tree partition $T^{\prime}$. Bags of $T^{\prime}$ that are not in $T$ have size at most $2$, and, to bags of $T^{\prime}$ that are also in $T$, we added at most $k^{2}$ vertices (at most one per edge between the bag and its parent). We conclude that $T^{\prime}$ has width at most $k(k+1)$.
A result of Ding and Oporowski [DO96] shows that tree-partition-width is tied to a parameter $\gamma$ that is (by design) monotonic with respect to the topological minor relation. We adapt their proof to derive the following stronger result.
@@
We use the following result of Chuzhoy and Tan [CT21] (the bound is weakened to have a lighter formula).
Lemma 5.10 (Chuzhoy and Tan [CT21]).
@@
Proof 5.12.
Let $G$ be a graph and $H$ be a topological minor of $G$. Any topological minor of $H$ is also a topological minor of $G$, hence $\operatorname{\textup{{wn}}}(G)\geq\operatorname{\textup{{wn}}}(H)$, $\gamma_{1}(G)\geq\gamma_{1}(H)$, $\gamma_{2}(G)\geq\gamma_{2}(H)$. We conclude that $\gamma(G)\geq\gamma(H)$.
@@
Since $C$ is connected, it contains a spanning tree $T$. We number its edges $e_{1},\dots,e_{\ell}$ such that every prefix induces a connected subtree of $T$. We construct a subgraph $H$ of $G$ that should be an $(m+1)$-multiple of a tree of order $m+1$, contradicting the definition of $m$. For each edge $uv$, in order, we try to add to $H$ $m+1$ vertex disjoint paths from $u$ to $v$ that avoid vertices of $C$ and the vertices already in $H$. If we manage to do this for at most $m$ edges, then we have placed at most $m(m+1)$ paths. Let $uv$ be the first edge for which we could not find $m+1$ vertex disjoint paths that do not intersect previous vertices (except for $u$ or $v$). By definition of $G^{b}$, there are $b$ vertex disjoint $u,v$-paths in $G$, we denote the set of such paths by $\pi$. At most $m$ of the paths of $\pi$ hit vertices of $C$ already in $H$. Then, since at least $b-m$ are hit by previous paths and there are at most $m(m+1)$ previous paths. By the pigeon hole principle, one of the previous paths $P_{0}$ must hit $\frac{b-m}{m(m+1)}\geq(m+1)^{2}$ paths in $\pi$. By considering $P_{0}$ and the paths it hits in $\pi$, we easily obtain a subdivision of an $(m+1)^{2}$-fan. This is a contradiction with the definition of $m$. Hence, we must have been able to process $m$ edges. Which means we obtained a subdivision of an $(m+1)$-multiple of a tree of order $m+1$. This is a contradiction to the definition of $m$. We conclude that the connected component must have size at most $m$.
@@
Proof 5.14.
Assume by contradiction that the maximum degree is more than $bm^{3}$. Let $B$ be a block of $H$, and $X$ be one if its vertices of maximum degree. $X$ contains at most $m$ vertices of $G$ by Claim 5. The vertices of $G$ in $B-X$ must be in the same connected component $C$ of $G$ since $b>m$. There are at least $bm^{3}+1$ edges between $X$ and $C$. By the pigeon hole principle, one vertex $v$ of $X$ must have at least $bm^{2}+1$ neighbours in $C$. Consider a spanning tree $T$ of $C$. We iteratively remove leaves that are not neighbours of $v$, and then replace any vertex of degree $2$ that is not a neighbour of $v$ by an edge between its neighbours. We denote this reduced tree by $T^{\prime}$.
First, note that the degree in $T^{\prime}$ is bounded by $b-1$ because incident edges can be extended to vertex disjoint paths to leaves of $T^{\prime}$ which are neighbours of $v$ by construction. We now use the fact that $G$ contains no $(m+1)$-branching-fans as topological minors. In particular, there must be a set $U$ of vertices of $T^{\prime}$ of size at most $m$ such that components of $T^{\prime}-U$ contain at most $m$ neighbours of $v$. By removing at most $m$ vertices of degree at most $(b-1)$, we have at most $1+(b-2)m$ components in $T^{\prime}-U$ meaning $v$ has degree bounded by $(b-1)m^{2}$. We found our contradiction.
Claim 7.
@@
Proof 5.15.
We first apply Lemma 5.10 to bound the treewidth of $G$ by $O(b)$. Consider a tree decomposition of $G$ of adequate width $\Theta(b)$, and replace each bag by the components of $G^{b}$ that intersect it. By Claim 2, this is a decomposition of $H$.
Using Claim 6 and Claim 7 and the construction of Wood as we did in the approximation algorithm, we obtain a tree-partition of $H$ of width $O(b^{2}m^{3})$. We then replace components of $G^{b}$ by their vertices, obtaining a tree-partition of $G$ of width $O(b^{2}m^{4})$ due to Claim 5.
We have obtained a tree-partition of width $O(b^{2}m^{4})=O(m^{24})$. This concludes the proof of Theorem 5.6.
@@ 6 Weighted Tree-Partition-Width
Recent investigations in algorithmic applications of tree-partition-width [BCvdW22, BMO+23] give algorithms that are fixed parameter tractable with the weighted tree-partition-width as parameter. In these cases, all vertices have weight one, and all edges have a positive weight. In this section, we show XALP-completeness for Weighted Tree-Partition-Width when all vertices and edges have weight one, and give an approximation algorithm similar to Theorem 3.11. The latter result can be regarded as a corollary of Theorem 3.11.
@@
Now, obtain a tree-partition of $G$ by undoing all contractions. I.e., if we contracted $vw$ to $x$ and $x\in B_{i}$, then replace $x$ by $v$ and $w$ in $B_{i}$. Repeat this step in the reverse order of which we did the contraction steps. The result is a tree-partition of $G$. Undoing contractions does not change the total weight of vertices in bags, so the total weight of vertices in a bag is bounded by $O(k^{7})$. Now, for each pair $ii^{\prime}$ of adjacent bags, there are $O(k^{14})$ pairs of vertices with one endpoint in each bag, and each edge between these bags has weight at most $k$, so the total weight of edges with one endpoint in $B_{i}$ and one endpoint in $B_{i^{\prime}}$ is bounded by $O(k^{15})$. Thus, the breadth of the obtained tree-partition is $O(k^{15})$.
@@
There is a polynomial time algorithm that constructs a tree-partition of breadth $O(k^{11}\log^{2}k)$ or reports that the weighted tree-partition-width is more than $k$.
There is an algorithm running in time $2^{O(k\log k)}n+2^{O(k)}n\log^{*}n$, that computes a tree-partition of breadth $O(k^{11})$ or reports that the weighted tree-partition-width is more than $k$.
Proof 6.4.
Use the same approach as in the previous algorithm, but use different subroutines to compute the approximate tree-partition, as in the different cases in Theorem 3.11.
As tree-partition-width is the special case of weighted tree-partition-width with all vertex weights one, and all edge weights zero, we directly have that deciding weighted tree-partition-width is XALP-hard. However, in applications, the case where all edges have positive weights is of main interest [BCvdW22, BMO+23]. In the following, we show that similar hardness also holds in some simple cases with edges having positive weights. We give a few intermediate results.
@@
Proof 6.6.
This can be proved in the same way as Lemma 4.1.
@@
Proof 6.8.
Take an instance of Tree-Partition-Width: $(G,k)$. Now, set the weight of all edges to 1, and all vertices to $L=k^{2}+1$. Now, for each tree-partition of $G$, its width (where we ignore weights) is at most $k$, if and only if the maximum over all bags of the total weight of vertices in a bag is at most $kL$, if and only if its breadth is at most $kL$. The latter follows as there are at most $k^{2}$ edges between bags each of weight one. The result now directly follows.
@@
Weighted Tree-Partition-Width with all vertex and edge weights 1 is XALP-complete.
Proof 6.10.
@@
Suppose the vertices in $C_{v,2}$ do not belong to the same bag. Then, by Claim 8, they belong to two adjacent bags, say $i_{1}$ and $i_{2}$. Now, $x_{v,1}$ and $x_{v,2}$ must be in bags, adjacent to $i_{1}$ and adjacent to $i_{2}$, so $x_{v,1},x_{v,2}\in B_{i_{1}}\cup B_{i_{2}}$. $B_{i_{1}}\cup B_{i_{2}}$ contain all $k+2$ vertices in $C_{v,2}\cup\{x_{v,1},x_{v,2}\}$; thus, within this set, there are at least $2k$ pairs of vertices with one endpoint in $B_{i_{1}}$ and one endpoint in $B_{i_{2}}$; however, only one pair of vertices in this set is not adjacent (namely, $x_{v,1}$ and $x_{v,2}$), so at least $12k-$ 1 edges cross the cut from $i_{1}$ to $i_{2}$, which contradicts the breadth of the tree-partition when $k>1$.
Claim 10.
@@
Proof 6.12.
Suppose $x_{v,3}\in B_{i}$. Now, as $x_{v,3}\cup C_{v,1}$ is a clique, all vertices in $x_{v,3}\cup C_{v,1}$ are in two incident bags, say $B_{i}\cup B_{i^{\prime}}$. As $B_{i}$ contains the $k$ vertices from $C_{v,2}$ (by Claim 9), $C_{v,1}\subseteq B_{i^{\prime}}$.
Claim 11.
@@
Proof 6.13.
Suppose the weighted tree-partition-width of $G^{\prime}$ is at most $k$. Suppose $(T,(B_{i})_{i\in V(T)})$ is a tree-partition of breadth at most $k$ of $G^{\prime}$. Let for all $i\in V(T)$, $B^{\prime}_{i}=\{v\in V\leavevmode\nobreak\ |\leavevmode\nobreak\ X_{v,1}\subseteq B_{i}\}$. It is easy to check that $(T,(B^{\prime}_{i})_{i\in V(T)})$ is a tree-partition of $G$ of breadth at most $k$. By Claim 10, each vertex $v\in V(G)$ belongs to one set $B^{\prime}_{i}$. As we replace a clique with $w_{V}(v)$ vertices of weight one by one vertex with weight $w_{V}(v)$, the total weight of each bag is still bounded by $k$. For each edge $\{v,w\}\in E(G)$, the bags containing $C_{v,1}$ and $C_{w,1}$ must be the same or adjacent, as there is an edge between a vertex in $C_{v,1}$ and a vertex in $C_{w,1}$. So, the weighted tree-partition-width of $G$ is at most $k$.
Now, suppose the weighted tree-partition-width of $G$ is at most $k$. Suppose $(T,(B_{i})_{i\in V(T)})$ is a tree-partition of breadth at most $k$ of $G$. Build a tree-partition of $G^{\prime}$ as follows: set $B^{\prime\prime}_{i}=\bigcup_{v\in B_{i}}C_{v,1}$, i.e., we replace each vertex $v$ by the set of vertices $C_{v,1}$, for all $i\in V(T)$. For each $v\in V(G)$, we add three extra bags: one bag $i_{v,2}$ containing all vertices in $C_{v,2}$, one bag containing only the vertex $x_{v,1}$ which is adjacent to bag $i_{v,2}$, and one bag containing only the vertex $x_{v,1}$ which also is adjacent to bag $i_{v,2}$. One easily checks that this is a tree-partition of $G^{\prime}$ of breadth at most $k$.
One can check that the transformation can be done with $O(k^{O(1)}+\log n)$ memory. So, we have a pl-reduction from an XALP-complete problem, and the result follows.

### v6

@@ 5 Tree-cut width and the stability of tree-partition-width
In this section, we consider the relation of the notion of tree-cut width with (stable) tree-partition-width. Tree-cut width was introduced by Wollan [32]. Ganian et al. [24] showed that several problems that are $W[1]$-hard with treewidth as parameter are fixed parameter tractable with tree-cut width as parameter.
We begin by defining the tree-cut width of a graph $G=(V,E)$. A tree-cut decomposition $(T,\mathcal{X})$ consists of a rooted tree $T$ and a family $\mathcal{X}$ of bags $(X_{i})_{i\in V(T)}$ which form a near partition of $V(G)$ (i.e. some bags may be empty, but nonempty bags form a partition of $V(G)$). For $t\in V(T)$, we denote by $e(t)$ the edge of $T$ incident to $t$ and its parent. For $e\in E(T)$, let $T_{1},T_{2}$ denote the two connected components of $T-e$. We denote by $\operatorname{cut}(e)$ the set of edges with an endpoint in both of $\bigcup_{i\in V(T_{1})}X_{i}$ and $\bigcup_{i\in V(T_{2})}X_{i}$. The adhesion of $t\in V(T)$ is $\text{adh}(t)=|\operatorname{cut}(e(t))|$, and its torso-size is $\text{tor}(t)=|X_{t}|+b_{t}$ where $b_{t}$ is the number of edges $e\in E(T)$ incident to $t$ such that $|\operatorname{cut}(e)|\geq 3$. The width of the decomposition is then $\max_{t\in V(T)}\{\text{adh}(t),\text{tor}(t)\}$. Note that edges are allowed to go between vertices that are not in the same bag. The tree-cut width of a graph is the minimal width of tree-cut decomposition. When $|\operatorname{cut}(e)|\geq 3$, the edge $e$ is called bold, and otherwise, $e$ is called thin. When $\text{adh}(t)\leq 2$, node $t$ is called thin, otherwise it is called bold. In [24], it is shown that a tree-cut decomposition can be assumed to be nice, meaning that if $t\in V(T)$ is thin then $N(Y_{t})\cap\left(\bigcup_{b\text{ sibling of }t}Y_{b}\right)=\varnothing$, where $Y_{i}$ is the union of $X_{j}$ for $j$ in the subtree of $i$.
Wollan [32] shows that having bounded tree-cut width is equivalent to only having wall immersions of bounded size.
Observation 5.1.
@@
It is easy to see that $K_{3,n}$ has a tree-cut decomposition of width $O(\sqrt{n})$: we place $A$ as the center of a star, with about $\sqrt{n}$ leaves of size $\sqrt{n}$. Now we consider an arbitrary tree-cut decomposition of $K_{3,n}$ achieving width $O(\sqrt{n})$. We first note that the vertices of $A$ cannot be split into separate bags because if they were, any edge of the decomposition on the path between such bags would have adhesion at least $n/2$. Hence, there is a bag containing $A$ and we may root the tree of our decomposition in this bag. Each subtree of the decomposition will contribute to the torso-size, and each vertex will contribute linearly to the adhesion of the edge from its subtree to the root. Since we assume the width to be $O(\sqrt{n})$, we must have at most $O(\sqrt{n})$ vertices per subtree. Consequently, there must be $\Omega(\sqrt{n})$ subtrees, so the torso-size of the root is $\Omega(\sqrt{n})$. ∎
@@
Proof.
Consider a nice tree-cut decomposition $(T,\mathcal{X})$ of a graph $G$ of width $k$. We will construct a tree-partition for a subdivision of $G$. Note that the bags are already disjoint, but some edges are not between neighboring bags of $T$.
@@
Suppose that $u,v$ are in the same bag of $T$ and the edge $uv$ was subdivided to form the path $u,a_{1},\dots,$ $a_{\ell},v.$ We add the vertices $a_{i}$ in new bags containing, $\{a_{1},a_{\ell}\}$, $\{a_{2},a_{\ell-1}\},\dots$ which corresponds to a new branch of the decomposition of width at most $2$.
@@
This gives a tree partition $T^{\prime}$. Bags of $T^{\prime}$ that are not in $T$ have size at most $2$, and, to bags of $T^{\prime}$ that are also in $T$, we added at most $k^{2}$ vertices (at most one per edge between the bag and its parent). We conclude that $T^{\prime}$ has width at most $k(k+1)$. ∎
A result of Ding and Oporowski [18] shows that tree-partition-width is tied to a parameter $\gamma$ that is (by design) monotonic with respect to the topological minor relation. We adapt their proof to derive the following stronger result.
@@
We use the following result of Chuzhoy and Tan [14] (the bound is weakened to have a lighter formula).
Lemma 5.8 (Chuzhoy and Tan [14]).
@@
Proof.
Let $G$ be a graph and $H$ be a topological minor of $G$. Any topological minor of $H$ is also a topological minor of $G$, hence $\operatorname{\textup{{wn}}}(G)\geq\operatorname{\textup{{wn}}}(H)$, $\gamma_{1}(G)\geq\gamma_{1}(H)$, $\gamma_{2}(G)\geq\gamma_{2}(H)$. We conclude that $\gamma(G)\geq\gamma(H)$. ∎
@@
Since $C$ is connected, it contains a spanning tree $T$. We number its edges $e_{1},\dots,e_{\ell}$ such that every prefix induces a connected subtree of $T$. We construct a subgraph $H$ of $G$ that should be an $(m+1)$-multiple of a tree of order $m+1$, contradicting the definition of $m$. For each edge $uv$, in order, we try to add to $H$ $m+1$ vertex disjoint paths from $u$ to $v$ that avoid vertices of $C$ and the vertices already in $H$. If we manage to do this for at most $m$ edges, then we have placed at most $m(m+1)$ paths. Let $uv$ be the first edge for which we could not find $m+1$ vertex disjoint paths that do not intersect previous vertices (except for $u$ or $v$). By definition of $G^{b}$, there are $b$ vertex disjoint $u,v$-paths in $G$, we denote the set of such paths by $\pi$. At most $m$ of the paths of $\pi$ hit vertices of $C$ already in $H$. Then, since at least $b-m$ are hit by previous paths and there are at most $m(m+1)$ previous paths. By the pigeon hole principle, one of the previous paths $P_{0}$ must hit $\frac{b-m}{m(m+1)}\geq(m+1)^{2}$ paths in $\pi$. By considering $P_{0}$ and the paths it hits in $\pi$, we easily obtain a subdivision of an $(m+1)^{2}$-fan. This is a contradiction with the definition of $m$. Hence, we must have been able to process $m$ edges. Which means we obtained a subdivision of an $(m+1)$-multiple of a tree of order $m+1$. This is a contradiction to the definition of $m$. We conclude that the connected component must have size at most $m$. ∎
@@
Proof.
Assume by contradiction that the maximum degree is more than $bm^{3}$. Let $B$ be a block of $H$, and $X$ be one if its vertices of maximum degree. $X$ contains at most $m$ vertices of $G$ by Claim 5.10. The vertices of $G$ in $B-X$ must be in the same connected component $C$ of $G$ since $b>m$. There are at least $bm^{3}+1$ edges between $X$ and $C$. By the pigeon hole principle, one vertex $v$ of $X$ must have at least $bm^{2}+1$ neighbors in $C$. Consider a spanning tree $T$ of $C$. We iteratively remove leaves that are not neighbors of $v$, and then replace any vertex of degree $2$ that is not a neighbor of $v$ by an edge between its neighbors. We denote this reduced tree by $T^{\prime}$.
First, note that the degree in $T^{\prime}$ is bounded by $b-1$ because incident edges can be extended to vertex disjoint paths to leaves of $T^{\prime}$ which are neighbors of $v$ by construction. We now use the fact that $G$ contains no $(m+1)$-branching-fans as topological minors. In particular, there must be a set $U$ of vertices of $T^{\prime}$ of size at most $m$ such that components of $T^{\prime}-U$ contain at most $m$ neighbors of $v$. By removing at most $m$ vertices of degree at most $(b-1)$, we have at most $1+(b-2)m$ components in $T^{\prime}-U$ meaning $v$ has degree bounded by $(b-1)m^{2}$. We found our contradiction. ∎
Claim 5.12.
@@
Proof.
We first apply Lemma 5.8 to bound the treewidth of $G$ by $O(b)$. Consider a tree decomposition of $G$ of adequate width $\Theta(b)$, and replace each bag by the components of $G^{b}$ that intersect it. By Claim 3.2, this is a decomposition of $H$. ∎
Using Claim 5.11 and Claim 5.12 and the construction of Wood as we did in the approximation algorithm, we obtain a tree-partition of $H$ of width $O(b^{2}m^{3})$. We then replace components of $G^{b}$ by their vertices, obtaining a tree-partition of $G$ of width $O(b^{2}m^{4})$ due to Claim 5.10.
We have obtained a tree-partition of width $O(b^{2}m^{4})=O(m^{24})$. This concludes the proof of Theorem 5.4. ∎
@@ 6 Weighted Tree-Partition-Width
Recent investigations in algorithmic applications of tree-partition-width [6, 11] give algorithms that are fixed parameter tractable with the weighted tree-partition-width as parameter. In these cases, all vertices have weight one, and all edges have a positive weight. In this section, we show XALP-completeness for Weighted Tree-Partition-Width when all vertices and edges have weight one, and give an approximation algorithm similar to Theorem 3.11. The latter result can be regarded as a corollary of Theorem 3.11.
@@
Now, obtain a tree-partition of $G$ by undoing all contractions. I.e., if we contracted $vw$ to $x$ and $x\in B_{i}$, then replace $x$ by $v$ and $w$ in $B_{i}$. Repeat this step in the reverse order of which we did the contraction steps. The result is a tree-partition of $G$. Undoing contractions does not change the total weight of vertices in bags, so the total weight of vertices in a bag is bounded by $O(k^{7})$. Now, for each pair $ii^{\prime}$ of adjacent bags, there are $O(k^{14})$ pairs of vertices with one endpoint in each bag, and each edge between these bags has weight at most $k$, so the total weight of edges with one endpoint in $B_{i}$ and one endpoint in $B_{i^{\prime}}$ is bounded by $O(k^{15})$. Thus, the breadth of the obtained tree-partition is $O(k^{15})$. ∎
@@
There is an algorithm running in polynomial time that constructs a tree-partition of breadth $O(k^{11}\log^{2}k)$ or reports that the weighted tree-partition-width is more than $k$.
There is an algorithm running in time $2^{O(k\log k)}n$, that computes a tree-partition of breadth $O(k^{11})$ or reports that the weighted tree-partition-width is more than $k$.
Proof.
Use the same approach as in the previous algorithm, but use different subroutines to compute the approximate tree-partition, as in the different cases in Theorem 3.11. ∎
As tree-partition-width is the special case of weighted tree-partition-width with all vertex weights one, and all edge weights zero, we directly have that deciding weighted tree-partition-width is XALP-hard. However, in applications, the case where all edges have positive weights is of main interest [6, 11]. In the following, we show that similar hardness also holds in some simple cases with edges having positive weights. We give a few intermediate results.
@@
Proof.
This can be proved in the same way as Lemma 4.1. ∎
@@
Proof.
Take an instance of Tree-Partition-Width: $(G,k)$. Now, set the weight of all edges to 1, and all vertices to $L=k^{2}+1$. Now, for each tree-partition of $G$, its width (where we ignore weights) is at most $k$, if and only if the maximum over all bags of the total weight of vertices in a bag is at most $kL$, if and only if its breadth is at most $kL$. The latter follows as there are at most $k^{2}$ edges between bags each of weight one. The result now directly follows. ∎
@@
Weighted Tree-Partition-Width with all vertex and edge weights equal to 1 is XALP-complete.
Proof.
@@
Suppose the vertices in $C_{v,2}$ do not belong to the same bag. Then, by Claim 6.6, they belong to two adjacent bags, say $i_{1}$ and $i_{2}$. Now, $x_{v,1}$ and $x_{v,2}$ must be in bags, adjacent to $i_{1}$ and adjacent to $i_{2}$, so $x_{v,1},x_{v,2}\in B_{i_{1}}\cup B_{i_{2}}$. $B_{i_{1}}\cup B_{i_{2}}$ contain all $k+2$ vertices in $C_{v,2}\cup\{x_{v,1},x_{v,2}\}$; thus, within this set, there are at least $2k$ pairs of vertices with one endpoint in $B_{i_{1}}$ and one endpoint in $B_{i_{2}}$; however, only one pair of vertices in this set is not adjacent (namely, $x_{v,1}$ and $x_{v,2}$), so at least $12k-$ 1 edges cross the cut from $i_{1}$ to $i_{2}$, which contradicts the breadth of the tree-partition when $k>1$. ∎
Claim 6.8.
@@
Proof.
Suppose $x_{v,3}\in B_{i}$. Now, as $x_{v,3}\cup C_{v,1}$ is a clique, all vertices in $x_{v,3}\cup C_{v,1}$ are in two incident bags, say $B_{i}\cup B_{i^{\prime}}$. As $B_{i}$ contains the $k$ vertices from $C_{v,2}$ (by Claim 6.7), $C_{v,1}\subseteq B_{i^{\prime}}$. ∎
Claim 6.9.
@@
Proof.
Suppose the weighted tree-partition-width of $G^{\prime}$ is at most $k$. Suppose $(T,(B_{i})_{i\in V(T)})$ is a tree-partition of breadth at most $k$ of $G^{\prime}$. Let for all $i\in V(T)$, $B^{\prime}_{i}=\{v\in V~{}|~{}X_{v,1}\subseteq B_{i}\}$. It is easy to check that $(T,(B^{\prime}_{i})_{i\in V(T)})$ is a tree-partition of $G$ of breadth at most $k$. By Claim 6.8, each vertex $v\in V(G)$ belongs to one set $B^{\prime}_{i}$. As we replace a clique with $w_{V}(v)$ vertices of weight one by one vertex with weight $w_{V}(v)$, the total weight of each bag is still bounded by $k$. For each edge $\{v,w\}\in E(G)$, the bags containing $C_{v,1}$ and $C_{w,1}$ must be the same or adjacent, as there is an edge between a vertex in $C_{v,1}$ and a vertex in $C_{w,1}$. So, the weighted tree-partition-width of $G$ is at most $k$.
Now, suppose the weighted tree-partition-width of $G$ is at most $k$. Suppose $(T,(B_{i})_{i\in V(T)})$ is a tree-partition of breadth at most $k$ of $G$. Build a tree-partition of $G^{\prime}$ as follows: set $B^{\prime\prime}_{i}=\bigcup_{v\in B_{i}}C_{v,1}$, i.e., we replace each vertex $v$ by the set of vertices $C_{v,1}$, for all $i\in V(T)$. For each $v\in V(G)$, we add three extra bags: one bag $i_{v,2}$ containing all vertices in $C_{v,2}$, one bag containing only the vertex $x_{v,1}$ which is adjacent to bag $i_{v,2}$, and one bag containing only the vertex $x_{v,1}$ which also is adjacent to bag $i_{v,2}$. One easily checks that this is a tree-partition of $G^{\prime}$ of breadth at most $k$. ∎
One can check that the transformation can be done with $O(k^{O(1)}+\log n)$ memory. So, we have a pl-reduction from an XALP-complete problem, and the result follows. ∎